A wire of length 16 inches is to be cut into two pieces; then each piece will be bent to form a square. Find the length of the two pieces assuming that the sum of the areas of the two squares is 10 square inches.
Answer by mananth(16946)

let one piece be x inches
other piece will be 16-x inches
..
(x/4)^2+(16-x)/4 ^2=10
x^2/16 +(256-32x+x^2)/16 =10
x^2+256-32x+x^2=160
2x^2-32x+96=0
/2
x^2-16x+48=0
(x-12)(x-4)=0
x=12 OR x=4
..
the pieces are 12 & 4 inches
...
m.ananth@hotmail.ca